tan square thita/(1+tan square thita). +. cot square thita/(1+cot square thita). =1

Grade:10

1 Answers

Yarra Deva
6 years ago
tan^2\Theta/(1+tan^2\Theta).+.cot^2\Theta/(1+cot^2\Theta)
put cot^2\Theta=1/tan^2\Theta (since \Theta not belongs to \pi/2 and 0)  (since cot and tan functions are undefined at 0 and \pi/2 )
=tan^2\Theta  /(1+tan^2\Theta  ) + (tan^2\Theta * 1/tan^2\Theta )/(1+tan^2\Theta )
=tan^2\Theta /(1+tan^2\Theta ) +  1/(1+tan^2\Theta )
=1